OpenStudy (anonymous):

If A (with arrow above)= (6.00i-8.00j) units, B (with arrow above)= (-8.00i+ 3.00j) units, and C (with arrow above) = (26.0i+ 19.0j) units, determine a and b such that aA(with arrow above+ bB(with arrow above) + C(with arrow above) = 0. a= b=

OpenStudy (anonymous):

?anyone

OpenStudy (anonymous):

6a + (-8)b + 26 = 0 -8a + 3b + 19 = 0

OpenStudy (anonymous):

Then you can multiply the top by 4 and the bottom by 3 24a - 32b + 104 = 0 -24a + 9b + 57 = 0 -23b + 161 = 0 161 = 23b b = 7 -8a + 3(7) + 19 = 0 -8a + 21 + 19 = 0 -8a + 40 = 0 40 = 8a 5 = a

OpenStudy (anonymous):

awesome!! thats what i got!!!! i just wanted to double check!!

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

No problem. Vector analysis is always fun!

OpenStudy (anonymous):

lol not for me only because she doesnt teach it correctly!